# Latent Action Pretraining From Videos

### v1

@@ 4.2 Baselines
For the choice of our underlying VLM , we utilize the 7B Large World Model (LWM-Chat-1M) (Liu et al., 2024), which also leads to much efficiency gains during pretraining (further analysis provided in Section 4.7). We also compare with existing behavior cloning baselines that do not require ground truth actions during pretraining.
@@
ActionVLA denotes the baseline that uses the actual ground-truth robot action labels during pretraining with the same backbone VLM. This may be seen as the upper bound, since it utilizes the actual ground-truth labels. However, we find that LAPA is more performant in scenarios where there are embodiment shifts between pretraining and fine-tuning (Section 4.5).
@@ 4.3 Language Table Results
Table 2: Language Table Results. Average Success Rate (%) across the three different pretrain-finetune combinations from the Language Table benchmark as described in Table 1. We also note the # of trajectories used for fine-tuning next to each category. We report the performance for individual tasks in Appendix E.1.
@@
First, we assess LAPA’s ability to learn from a small subset of in-domain action label data by pretraining on 181k trajectories and finetuning on 1k action-labeled trajectories (0.5%). As shown in Table 2, LAPA largely outperforms Scratch and narrows the gap with ActionVLA despite not using action labels during pretraining. Additionally, LAPA surpasses UniPi and Vpt. Notably, while UniPi handles simple tasks well, its diffusion model often generates incorrect plans for longer-horizon tasks, aligning with Du et al. (2024) (see Figure 17 of Appendix E.1). Vpt, with the same backbone VLM as LAPA, outperforms UniPi, showing the superiority of the VLA model, but still underperforms LAPA, highlighting the effectiveness of latent actions.
@@
We investigate whether LAPA’s broad skills can be retained after finetuning on a specific task. Pretraining LAPA on 181k trajectories and finetuning on only separate tasks (7k), we evaluate all 5 task categories, similar to the in-domain setup, to assess latent pretraining’s benefits for unseen tasks. When comparing LAPA and Scratch in Table 2 and Table 6, 7 in Appendix E.1, latent pretraining significantly benefits the separate task as well the other 4 task categories, resulting in a significant boost in both seen and unseen setups. Like before, UniPi is constrained by its diffusion model’s planning limitations, while Vpt performs strongly, even surpassing ActionVLA in the unseen setting. This is likely due to using more labeled data (7k vs. 1k), helping the IDM generate more accurate pseudo labels.
@@
We further investigate if LAPA benefits downstream performance when the pretraining and fine-tuning environments are different. We pretrain LAPA on 440k real-world trajectories, and then finetune on 1k simulation trajectories, which can be seen as testing on a setup where a real2sim gap is present.  [^5] We only test on the real2sim scenario since simulation evaluation is easily replicable and easier to setup. We leave exploration in the sim2real scenario as future work. Figure 3 (a) compares the two environments. From Table 2, we observe that LAPA still significantly outperforms Scratch, showing that latent pretraining leads to positive transfer even on cross-environment setting. Notably, both UniPi and Vpt significantly underperforms LAPA, showing that learning to predict latent actions is more robust to cross-environment transfer. Vpt only results in minor positive transfer, indicating that the IDM is not robust to environment shifts.
4.4 SIMPLER Results
We pretrain our models on the Bridgev2 (Walke et al., 2023) dataset and fine-tune on 100 trajectories collected from the SIMPLER environment (Li et al., 2024b) (25 trajectories for each task).
Figure 4: SIMPLER Results. Avg. success rate (%) is shown across 4 tasks. Detailed results are in Appendix E.2.
Results
As shown in Figure 4, UniPi significantly underperforms all other baselines on the SIMPLER Environment. We observe that, although the generated plans from the diffusion models are quite accurate, the IDM lacks the capability to predict 7 DOF continuous actions accurately when given only 100 action-labeled trajectories. Specifically, we observe that UniPi often fails to grasp the object within the maximum step limit. This implies the effectivness of using VLAs in scenarios with insufficient action-labeled data. Similar to the results of Section 4.3, LAPA outperforms baseline models that pretrain on actionless videos (UniPi and Vpt) and closes the performance gap with ActionVLA, which is pretrained on all of the 60K action-labeled trajectories from the Bridgev2 dataset. This highlights the effectiveness of LAPA, even when the complexity of the action space increases. We provide the detailed results of individual tasks and subtasks (grasping and reaching) in Table 10 of Appendix E.2.
@@ 4.5 Real-world Results
Figure 5: Real-world Tabletop Manipulation Results. We evaluate on a total of 54 rollouts for each model encompassing unseen object combinations, unseen objects and unseen instructions. Average success rate (%) are shown. We provide detailed results depedning on the generalization type in Table 12 and individual results in Appendix E.3.
We pretrain our models on (1) Bridgev2 (Walke et al., 2023) to measure the cross-embodiment performance (WidowX embodiment for pretraining and Franka embodiment for finetuning) and (2) Open X-Embodiment Dataset  (Collaboration et al., 2023) to measure the effect of pretraining in a multi-embodiment setting. Figure 5 shows the average success rate across the 3 tasks where each task encompasses unseen object combination, unseen object, and unseen instruction settings. We provide detailed results depending on the generalization type in Table 12 in Appendix E.3. We also measure the cross-embodiment performance of Open X-Embodiment Dataset pretraining by fine-tuning on bi-manual robots.
@@
We compare models that were pretrained on the Bridgev2 dataset. Similar to previous results, all models pretrained on Bridgev2 result in significant performance enhancement compared to Scratch. Furthermore, by comparing LAPA which does not leverage action-labeled trajectories during pretraining with models that use action-labeled trajectories during pretraining (ActionVLA and OpenVLA (Bridge)), we observe an interesting finding: LAPA outperform VLAs that use action labeled pretraining data on average success rate of the 3 tasks, unlike previous scenarios where VLAs pretrained on the ground-truth actions were upper bounds. LAPA significantly outperforms the other models in pick-and-place tasks; given that most tasks in Bridgev2 are pick-and-place, we hypothesize that VLA models pretrained on ground truth action labels have overfitted to the WidowX action space from the Bridgev2 dataset, hampering cross-embodiment adaptability to action distribution shifts during fine-tuning. In contrast, LAPA avoids this issue by not relying on ground truth action labels during pretraining.
@@
When comparing LAPA with OpenVLA, we see that LAPA significantly outperforms OpenVLA on 2 out of 3 tasks (Figure 5). This highlights LAPA’s effectiveness in a multi-embodiment setting by showcasing its ability to leverage a shared latent action space during pretraining, akin to how language and image representations are utilized. In contrast, contemporary action pretraining methods may suffer from reduced positive transfer between datasets due to the variability in action representation spaces across different embodiments and datasets.
However, for pick and place task, LAPA underperforms OpenVLA. We observe that most failures of LAPA are due to early grasping. In fact, LAPA outperforms OpenVLA in reaching performance (83.33% vs 66.67%) (reaching performance for each task is provided separately in Appendix E.3). This suggests that, although LAPA possesses stronger language conditioning and coarse-grained planning abilities, there is room for improvement in skills such as grasping. Since grasping occurs only once or twice in each trajectory, the 150 labeled trajectories may not be sufficient for LAPA to accurately predict grasp actions based on the physical characteristics of diverse objects.
To evaluate the cross-embodiment performance of LAPA pretrained with Open-X, we fine-tune both LAPA and OpenVLA on a bimanual robot with a 14-DoF action space, which presents a more challenging cross-embodiment scenario than the previous Bridgev2 pretraining experiments due to a cross-embodiment gap as well as the increase of the action space dimensions (there are no bimanual robot datasets in Open-X). Similar to the single-arm experiments, our evaluation covers seen object combinations, unseen object combinations, unseen objects, and unseen instructions. We observe that LAPA (30.21%), which does not use labeled action data during pretraining, slightly outperforms OpenVLA (26.04%) on average, confirming LAPA’s cross-embodiment capability even with the increase of the complexity of the action dimensions. However, the absolute performance of both models remains relatively low, indicating much room for improvement. We believe that incorporating bimanual human manipulation videos (Grauman et al., 2022; Damen et al., 2022) could enhance LAPA’s performance, which we plan to explore in future work. Individual results are provided in Table 17 in Appendix E.3.
@@ 4.6 Learning from Human Manipulation Videos
In this section, we show results when we extend LAPA to human manipulation videos, which aligns with the main motivation of this work. Unlike robot trajectories, human videos have two challenges: human videos do not contain action labels, and the distribution of human videos is distinct from the robot embodiment (McCarthy et al., 2024). We try to investigate whether our method as well as baseline approaches could address these challenges by pretraining on Something-Something V2 dataset (Goyal et al., 2017) which consists of 220K videos that includes human performing actions with everyday objects.
@@
Figure 6: Pretraining from Human Video Results. Average success rate (%) of LAPA and baselines pretrained on human manipulation videos where the embodiment and environment gap is extreme. We evaluate on both simulation (left) and real-world robot setup (right).
We first evaluate the performance of LAPA pretrained on human videos on SIMPLER. In addition to Scratch, we also compare with UniPi and Vpt pretrained with the same human video dataset. As shown in Figure 6(a), LAPA outperforms Scratch, showing that although the distribution of the pretraining data is distinct from the deployment setup, leveraging human videos for latent action pretraining results in positive transfer. Also, consistent with the result of Section 4.4, LAPA shows the best performance, implying that Latent Action Pretraining is robust to human to robot embodiment shifts. Note that it is impossible to train ActionVLA because the human videos do not have any robot action labels.
We report the real-world robot experiments in Figure 6(b). Surprisingly, we can see that LAPA trained with human videos outperforms OpenVLA (Bridge) on average. Despite the larger embodiment gap for LAPA (Human to robot vs. Robot to robot), it learns a better prior for robot manipulation. This result highlights the potential of raw human manipulation videos from the web compared to expensive robot manipulation data, which requires time-intensive teleoperation to collect. We expect that applying our approach on large-scale internet videos (e.g., YouTube videos) could unlock the potential for large-scale pretraining of a generalist action foundational model, similar to foundational models in Natural Language Processing or Computer Vision.
@@ 5.1 Scaling Model, Data, and Latent Action Size
(b) Data Scaling (%)
(c) Latent Action Seq
@@
Large Language Models (LLMs) have demonstrated scaling laws (Kaplan et al., 2020), where performance improves with increases in model size, dataset size, and computational resources used for training. Similarly, we attempt to analyze whether LAPA benefits from scaling across four dimensions: latent action quantization model size, data size, latent action sequence and vocabulary size. For a controlled setup, we apply our method to Bridgev2 and then fine-tune it on SIMPLER.
As shown in Figure 7, scaling benefits LAPA across the four dimensions. However, for the latent action space (sequence and vocabulary size), we believe there is a sweet spot where performance improves up to a certain point, after which it either plateaus or diminishes. Interestingly, we observe that the optimal scale of the latent action space depends on the complexity of the action dimension contained in the pretraining dataset. For example, increasing the latent action sequence length is less effective compared to increasing the vocabulary for Language Table (Figure 16). Except for Language Table, we maintain the generation space of LAPA at $8^{4}$ throughout all of our main experiments.
These results imply that when scaling pretraining to Internet-scale videos that go beyond manipulation videos, scaling LAPA in terms of model, dataset, and latent action space could improve performance, especially to capture higher action dimensions such as whole-body locomotion and manipulation.
@@ 5.2 Latent Action Analysis
Figure 8: Latent Action Analysis in Language Table. We condition the current observation $x_{1}$ and quantized latent action to the decoder of the latent action quantization model. We observe that each latent action can be mapped into a semantic action. For example, latent action 0 corresponds to moving a bit left and forward.
Figure 9: Correlation of latent actions with ground-truth actions.
First, we analyze latent actions learned from Language Table with vocabulary size of 8 and sequence length of 1. In Figure 8, we show that each latent action corresponds to a distinct movement of the robot arm, with the distribution of latent actions being well-clustered in the actual 2D action space. We observe that increasing the latent action vocabulary size leads to capturing a more fine-grained information. We analyze the relationship between latent actions with ground-truth 2 DOF actions by mapping each instance into latent action space. As shown in Figure 9, we observe that latent actions are well-clustered in the actual 2D action space, indicating that latent actions are meaningful representations that are highly related to actual continuous actions.
We further analyze the latent actions learned from human manipulation videos using the Something-Something V2 dataset. As illustrated in Figure 10, these latent actions capture not only hand movements but also camera movements. Since the camera viewpoint varies throughout the videos in the Something-Something V2 dataset due to the videos being egocentric, our latent action quantization model also learns to represent camera movements.
Figure 11: Latent Action Analysis in Multi-Embodiment Setting. We condition the current observation $x_{1}$ and quantized latent action to the decoder of the latent action quantization model. We observe that each latent action can be mapped into a similar semantic action even though the embodiments are different. For example, latent action [1,1,3,2] corresponds to going down and left while [3,2,0,1] corresponds to going up a little bit.
We also analyze the latent actions learned from the Open-X embodiment, which encompasses multiple embodiments, tasks, and environments. As shown in Figure 11, even though the embodiment and environment differ, conditioning on the same latent action results in a similar action in the reconstructed image. This supports our previous claim that latent actions are learned in a shared representation space, regardless of the embodiment or dataset, facilitating stronger positive transfer across diverse datasets.
@@
We also qualitatively analyze the coarse-grained planning capability of LAPA through a closed-loop rollout. We use a LAPA model that has only undergone pretraining, without any action finetuning. Since this model generates latent actions that are not directly executable in the real world, we condition the current observation $x_{1}$ and the predicted latent action from LAPA with the decoder of the latent action quantization model. As shown in Figure 12, when conditioned on the current observation and the instruction to ‘take the broccoli out of the pot’, LAPA generates robot trajectories that successfully reaches for the broccoli, moves down to grab it, and, as the arm moves away from the pot, the broccoli disappears. This shows the potential for LAPA as a general-purpose robotic world model, not only predicting actions but also the outcomes of the actions. For example, this can lead to an extension of LAPA to act as a Task and Motion planning system, where it can first generate multiple plans given a natural language task instruction, choose the most optimal trajectory based on methods of quantifying the success among multiple trajectory candidates (Hwang et al., 2024; Duan et al., 2024), and perform open-loop / closed-loop inference. This can lead a paradigm where we aim to improve end-task performance through scaling test-time compute, as with LLMs (Snell et al., 2024).

### v2

@@ 4.2 Baselines
For the underlying VLM, we use the 7B Large World Model (LWM-Chat-1M) (Liu et al., 2024).
@@
ActionVLA denotes the baseline that uses the actual ground-truth robot action labels during pretraining with the same backbone VLM. This may be seen as the upper bound, since it utilizes the actual ground-truth labels.
@@ 4.3 Language Table Results
Table 1: Language Table Results. Average Success Rate (%) $\pm$ StdErr across the three different pretrain-finetune combinations from the Language Table benchmark as described in Table 3. We also note the # of trajectories used for fine-tuning next to each category.
@@
First, we assess LAPA’s ability to learn from a small subset of in-domain action label data by pretraining on 181k trajectories and finetuning on 1k action-labeled trajectories (0.5%). As shown in Table 1, LAPA largely outperforms Scratch and narrows the gap with ActionVLA despite not using action labels during pretraining. Additionally, LAPA surpasses UniPi and Vpt. Notably, while UniPi handles simple tasks well, its diffusion model often generates incorrect plans for longer-horizon tasks, aligning with Du et al. (2024) (see Figure 17 of Appendix G.1). Vpt, with the same backbone VLM as LAPA, outperforms UniPi, showing the superiority of the VLA model, but still underperforms LAPA, highlighting the effectiveness of latent actions.
@@
We investigate whether LAPA’s broad skills can be retained after finetuning on a specific task. Pretraining LAPA on 181k trajectories and finetuning on only separate tasks (7k), we evaluate all 5 task categories, similar to the in-domain setup, to assess latent pretraining’s benefits for unseen tasks. When comparing LAPA and Scratch in Table 1 and Table 7, 8 in Appendix G.1, latent pretraining significantly benefits the separate task as well the other 4 task categories, resulting in a significant boost in both seen and unseen setups. Like before, UniPi is constrained by its diffusion model’s planning limitations, while Vpt performs strongly, even surpassing ActionVLA in the unseen setting. This is likely due to using more labeled data (7k vs. 1k), helping the IDM generate more accurate pseudo labels.
@@
We further investigate if Latent Action Pretraining benefits downstream performance when the pretraining and fine-tuning environments are different. We pretrain LAPA on 440k real-world trajectories, and then finetune on 1k simulation trajectories, which can be seen as testing on a setup where a real2sim gap is present (Figure 9 (a)). From Table 1, we observe that LAPA still significantly outperforms Scratch, showing that latent pretraining leads to positive transfer even on cross-environment setting. Notably, both UniPi and Vpt significantly underperforms LAPA, showing that learning to predict latent actions is more robust to cross-environment transfer. Vpt only results in minor positive transfer, indicating that the IDM is not robust to environment shifts.
@@ 4.4 Real-world Results
We pretrain our models on (1) Bridgev2 (Walke et al., 2023) to measure the cross-embodiment performance (WidowX embodiment for pretraining and Franka embodiment for finetuning) and (2) Open X-Embodiment Dataset  (Collaboration et al., 2023) to measure the effect of pretraining in a multi-embodiment setting. Figure 3 shows the average success rate across the 3 tasks where each task encompasses unseen object combination, unseen object, and unseen instruction settings. We provide detailed results depending on the generalization type in Table 2.
Figure 3: Real-world Tabletop Manipulation Results. We evaluate on a total of 54 rollouts for each model encompassing unseen object combinations, unseen objects and unseen instructions. Average success rate (%) $\pm$ StdErr are shown (detailed results provided in Appendix G.3).
@@
We compare models that were pretrained on the Bridgev2 dataset. Similar to previous results, all models pretrained on Bridgev2 result in significant performance enhancement compared to Scratch. Furthermore, by comparing LAPA which does not leverage action-labeled trajectories during pretraining with models that use action-labeled trajectories during pretraining (ActionVLA and OpenVLA), we observe an interesting finding: LAPA outperform VLAs that use action labeled pretraining data on average success rate of the 3 tasks, unlike previous scenarios where VLAs pretrained on the ground-truth actions were upper bounds. LAPA significantly outperforms the other models in pick-and-place tasks; given that most tasks in Bridgev2 are pick-and-place, we hypothesize that VLA models pretrained on ground truth action labels have overfitted to the WidowX action space from the Bridgev2 dataset, hampering cross-embodiment adaptability to action distribution shifts during fine-tuning. In contrast, LAPA avoids this issue by not relying on ground truth action labels during pretraining.
Seen Obj. Unseen Seen Obj. AVG
Unseen Combo Obj. Unseen Instr.
Scratch 18.0 20.3 25.4 21.2
ActionVLA (Bridge) 38.3 31.8 27.7 32.6
OpenVLA (Bridge) 35.6 34.6 22.1 30.8
LAPA (Bridge) 43.4 31.4 35.6 36.8
OpenVLA (Open-X) 46.2 42.1 43.4 43.9
LAPA (Open-X) 57.8 43.9 48.5 50.1
LAPA (Human Videos) 36.5 37.4 28.1 34.0
@@
When comparing LAPA with OpenVLA, we see that LAPA significantly outperforms OpenVLA on 2 out of 3 tasks (Figure 3). Also, as shown in Table 2, LAPA (Open-X) outperforms OpenVLA (Open-X) on all types of generalization settings. This highlights LAPA’s effectiveness in a multi-embodiment setting by showcasing its ability to leverage a shared latent action space during pretraining, akin to how language and image representations are utilized. In contrast, contemporary action pretraining methods may suffer from reduced positive transfer between datasets due to the variability in action representation spaces across different embodiments and datasets.
However, for pick and place task, LAPA underperforms OpenVLA. We observe that most failures of LAPA are due to early grasping. In fact, LAPA outperforms OpenVLA in reaching performance (83.33% vs 66.67%) (reaching performance for each task is provided separately in Appendix G.3). This suggests that, although LAPA possesses stronger language conditioning and coarse-grained planning abilities, there is room for improvement in skills such as grasping. Since grasping occurs only once or twice in each trajectory, the 150 labeled trajectories may not be sufficient for LAPA to accurately predict grasp actions based on the physical characteristics of diverse objects.
@@ 4.5 Learning from Human Manipulation Videos
Figure 4: Pretraining from Human Video Results. Average success rate (%) $\pm$ StdErr of LAPA and baselines pretrained on human manipulation videos where the embodiment and environment gap is extreme. We evaluate on both simulation (left) and real-world robot setup (right).
In this section, we show results when we extend Latent Action Pretraining to human manipulation videos, which aligns with the main motivation of this work. Unlike robot trajectories, human videos have two challenges: human videos do not contain action labels, and the distribution of human videos is distinct from the robot embodiment (McCarthy et al., 2024). We try to investigate whether our method as well as baseline approaches could address these challenges by pretraining on Something-Something V2 dataset (Goyal et al., 2017) which consists of 220K videos that includes human performing actions with everyday objects.
We first evaluate the performance of LAPA pretrained on human videos on SIMPLER. In addition to Scratch, we also compare with UniPi and Vpt pretrained with the same human video dataset. As shown in Figure 4(a), LAPA outperforms Scratch, showing that although the distribution of the pretraining data is distinct from the deployment setup, leveraging human videos for latent action pretraining results in positive transfer. Also, LAPA performs the best performance by outperforming UniPi and Vpt, implying that Latent Action Pretraining is robust to human to robot embodiment shifts. Note that it is impossible to train ActionVLA because the human videos do not have any robot action labels.
We report the real-world robot experiments in Figure 4(b). Surprisingly, we can see that LAPA trained with human videos outperforms OpenVLA (Bridge) on average. Despite the larger embodiment gap for LAPA (Human to robot vs. Robot to robot), it learns a better prior for robot manipulation. Also, as shown in Table 2, LAPA (Human Videos) shows good generalization performance, especially for unseen objects. We conjecture that this is because Something Something V2 dataset interacts with much diverse objects compared to Bridgev2. This result highlights the potential of raw human manipulation videos from the web compared to expensive robot manipulation data, which requires time-intensive teleoperation to collect. We expect that applying our approach on large-scale internet videos (e.g., YouTube videos) could unlock the potential for large-scale pretraining of a generalist action foundational model, similar to foundational models in NLP or Computer Vision.
@@ 5.1 Scaling Model, Data, and Latent Action Size
(b) Data Scaling
(c) Latent Action Length
@@
Large Language Models (LLMs) have demonstrated scaling laws (Kaplan et al., 2020), where performance improves with increases in model size, dataset size, and computational resources used for training. Similarly, we attempt to analyze whether LAPA benefits from scaling across three dimensions: latent action quantization model size, data size, and latent action representation space. For a controlled setup, we apply our method to Bridgev2 and then fine-tune it on SIMPLER except for Language Table result of Figure 5(c).
As shown in Figure 5, scaling benefits LAPA across the three dimensions. Interestingly, we observe that the optimal scale of the latent action space depends on the complexity of the action dimension contained in the pretraining dataset. For example, increasing the latent action sequence length is less effective compared to increasing the vocabulary for Language Table (Figure 16). Except for Language Table, we maintain the generation space of LAPA at $8^{4}$ throughout all of our main experiments. These results imply that when scaling pretraining to Internet-scale videos that go beyond manipulation videos, scaling LAPA in terms of model, dataset, and latent action space could improve performance, especially to capture higher action dimensions such as whole-body control.
@@ 5.2 Latent Action Analysis
Figure 6: Latent Action Analysis. We condition the current observation $x_{1}$ and quantized latent action to the decoder of the latent action quantization model. We observe that each latent action can be mapped into a semantic action. For example, latent action [1,1,3,2] corresponds to going down and left while [3,2,0,1] corresponds to going up a little bit.
@@
In Language Table, we observe that each latent action corresponds to a distinct movement of the robot arm, with the distribution of latent actions being well-clustered in the actual 2D action space (shown in Figure 12, 13 of Appendix E). Next, for human manipulation videos, we observe that camera viewpoints also correspond to a latent action since the viewpoint changes within a video (shown in Figure 14 of Appendix E). We also analyze the latent actions learned from the Open-X embodiment, which encompasses multiple embodiments, tasks, and environments. As shown in Figure 6, even though the embodiment and environment differ, conditioning on the same latent action results in a similar action in the reconstructed image. This supports our previous claim that latent actions are learned in a shared representation space, regardless of the embodiment or dataset, facilitating stronger positive transfer across diverse datasets.
@@
We also qualitatively analyze the coarse-grained planning capability of LAPA through a closed-loop rollout. We use a LAPA model that has only undergone pretraining, without any action finetuning. Since this model generates latent actions that are not directly executable in the real world, we condition the current observation $x_{1}$ and the predicted latent action from LAPA with the decoder of the latent action quantization model. As shown in Figure 7, when conditioned on the current observation and the instruction to ‘take the broccoli out of the pot’, LAPA generates robot trajectories that successfully reaches for the broccoli, moves down to grab it, and, as the arm moves away from the pot, the broccoli disappears. This shows the potential for LAPA as a general-purpose robotic world model, not only predicting actions but also the outcomes of the actions. For example, this can lead to an extension of LAPA to act as a Task and Motion planning system, where it can first generate multiple plans given a natural language task instruction, choose the most optimal trajectory based on methods of quantifying the success among multiple trajectory candidates (Hwang et al., 2024; Duan et al., 2024), and perform open-loop / closed-loop inference. This can lead a paradigm where we aim to improve performance through scaling test-time compute, as with LLMs (Snell et al., 2024).